# Singular-degenerate parabolic systems with the conormal boundary condition on the upper half space

### v1

@@ Abstract.
We prove the well-posedness and regularity of solutions in mixed-norm weighted Sobolev spaces for a class of second-order parabolic and elliptic systems in divergence form in the half-space $\mathbb{R}^{d}_{+}=\{x_{d}>0\}$ subject to the conormal boundary condition. Our work extends results previously available for scalar equations to the case of systems of equations. The leading coefficients are the product of $x_{d}^{\alpha}$ and bounded non-degenerate matrices, where $\alpha\in(-1,\infty)$. The leading coefficients are assumed to be merely measurable in the $x_{d}$ variable, and to have small mean oscillations in small cylinders with respect to the other variables. Our results hold for systems with lower-order terms that may blow-up near the boundary.
@@ 1. Introduction
Here for $i,j=1,\dots,d$, the coefficients $A^{ij}$, $B^{i}$, $\hat{B}^{i}$, $C$, $C_{0}:\Omega_{T}\to\mathbb{R}^{n\times n}$ and $a_{0}=a_{0}(x_{d}):\mathbb{R}_{+}\to\mathbb{R}^{n\times n}$ are $n\times n$ matrix-valued measurable functions on $\Omega_{T}$, $F^{i},F_{0},f:\Omega_{T}\to\mathbb{R}^{n}$ are given (column-) vector-valued measurable functions on $\Omega_{T}$, $u:\Omega_{T}\to\mathbb{R}^{n}$ is the unknown solution, and $\lambda\geq 0$ is a parameter. We impose standard non-degeneracy, boundedness, and strong ellipticity conditions on $a_{0},C_{0}$, and $A^{ij}$. See (2.1) and (2.2) in Section 2.
@@
We show that under some regularity assumptions on $A^{ij}$ and $C_{0}$, for any $p,q\in(1,\infty)$ and wide classes of weights $w=w(t,x)=w_{0}(t)w_{1}(x)$, the estimate
$\|Du\|_{L_{q,p}(\Omega_{T},wd\mu)}+\sqrt{\lambda}\|u\|_{L_{q,p}(\Omega_{T},wd\mu)}\leq N(\|F\|_{L_{q,p}(\Omega_{T},wd\mu)}+\|f\|_{L_{q,p}(\Omega_{T},wd\mu)})$ (1.4)
holds, where $N>0$ does not depend on $u,F$, and $f$, $\lambda>0$ is sufficiently large, and $L_{q,p}(\Omega_{T},wd\mu)$ is a mixed-norm weighted Lebesgue space with the norm
$\|f\|_{L_{q,p}(\Omega_{T},wd\mu)}=\left(\int_{-\infty}^{T}\left(\int_{\mathbb{R}^{d}_{+}}|f(t,x)|^{p}w_{1}(x)x_{d}^{\alpha}dx\right)^{q/p}w_{0}(t)dt\right)^{1/q}.$
Here $d\mu$ stands for $x_{d}^{\alpha}dxdt$, which is a natural measure for $L_{p}$ estimates for (1.1)-(1.2). In addition to estimate (1.4), we also obtain the corresponding existence and uniqueness of solutions. See Theorem 2.2 (i) and Section 2 for precise statements and definitions. In the case of scalar equations ($n=1$), the result of our Theorem 2.2 (i) was known from the recent work [17]. The arguments in [17] cannot be directly generalized to the case of systems due to the use of Moser’s iteration, a purely scalar technique, in one of the key steps of the proof. Showing that estimate (1.4) and solvability of (1.1)-(1.2) still hold in the case of systems with $n\geq 2$ was one of our motivations to write this paper.
@@
In [16] and [17] the authors considered scalar ($n=1$) non-divergence form analogues of (1.1)-(1.2) in the ranges $\alpha\in(-1,1)$ and $\alpha\in(-1,\infty)$, respectively. There, an additional structural condition $A^{dj}/A^{dd}\equiv\text{const}$ for $1\leq j\leq d-1$ is imposed. The results in [16, 17] essentially use the scalar structure of the equations and cannot be generalized to systems of equations. Moreover, it remains an open problem whether classical $W^{2}_{p}$ estimates hold for elliptic and parabolic systems with partially VMO coefficients. We therefore do not consider systems in non-divergence form in this paper and leave this question for future investigation.
@@
Recent results on boundary Schauder-type estimates for parabolic equations with the conormal boundary condition can be found in [1, 2, 9], and for elliptic equations in [10, 42]. In these works, Schauder estimates for equations with singular-degenerate coefficients are used to derive boundary Harnack principles for equations with bounded uniformly elliptic coefficients. Parabolic and elliptic systems in divergence form with the zero Dirichlet boundary condition were studied in [14], whereas scalar equations in non-divergence form were considered in [18]. In our recent work [3], we studied the solvability of equations in divergence form with the inhomogeneous Dirichlet boundary data taken in spaces of distributions with negative order of differentiability. In these papers, the natural range of the parameter $\alpha$ for equations with the Dirichlet boundary condition is $\alpha\in(-\infty,1)$.
@@
Various phenomena in pure and applied mathematics are modeled by equations whose coefficients in front of $u_{t},D^{2}u,Du$, and $u$ may exhibit different rates of decay or blow-up. Examples of such anisotropy in coefficients include the Black-Scholes-Merton equation in mathematical finance, degenerate viscous Hamilton-Jacobi equations, and equations involving Baouendi-Grushin operators arising in sub-Riemannian geometry. For recent results on weighted $L_{p}$ estimates for such anisotropic equations we refer the reader to [23, 22, 20, 21, 39].
@@
Lastly, $L_{p}$ estimates for elliptic and parabolic equations with unbounded lower-order terms is an extensively studied topic, particularly in the classical case $\alpha=0$. Such results are often obtained under some integrability assumptions on the coefficients $B,\hat{B}$, and $C$; common conditions include $B,\hat{B}\in L_{d}$, $C\in L_{d/2}$, or their generalizations. Under these assumptions, $L_{p}$ estimates can be established for both Dirichlet and conormal boundary conditions. For detailed results in this direction, see [28, 38, 34] for elliptic equations and [33] for parabolic equations. As we mentioned earlier, $L_{p}$ estimates for equations with the zero Dirichlet boundary condition under assumptions of type (1.5) are also well-studied in the literature; see [12, 30, 35].
@@
The rest of the paper is organized as follows. In Section 2 we introduce the notation, function spaces, and state the main results, Theorems 2.2 and 2.3. In Section 3 we prove the existence and uniqueness of $L_{2}$ solutions. In Section 4 we consider systems having simple coefficients. In Section 5 we prove Part (i) of Theorem 2.2, where $B,\hat{B}$, and $C$ satisfy (1.3). Finally, Section 6 addresses systems with singular lower-order terms satisfying (1.5), completing the proofs of Theorems 2.2 and 2.3.
@@ 2.1. Notation
By $D_{i}u=\frac{\partial u}{\partial x_{i}}$, $i=1,\dots,d$, we denote the spatial weak partial derivatives of $u$, $Du=(D_{1}u,\dots,D_{d}u)$, and $D_{x^{\prime}}u=(D_{1}u,\dots,D_{d-1}u)$. We set $|A|^{2}=\sum_{a\in A}|a|^{2}$, where $A$ is a scalar, vector, matrix, or their collection, and the sum is taken over all scalar entries $a$ of $A$.
@@
$\kappa\sum_{i=1}^{d}|\xi_{i}|^{2}\leq\sum_{i,j=1}^{d}A^{ij}\xi_{j}\cdot\xi_{i},\quad|A^{ij}|\leq\kappa^{-1},$ (2.1)
on $\Omega_{T}$ for all $\xi_{i}\in\mathbb{R}^{n},i=1,\dots,d$. Here $a\cdot b:=\sum_{k=1}^{n}a^{k}b^{k}$ denotes the dot product in $\mathbb{R}^{n}$. The coefficient matrices $a_{0}$ and $C_{0}$ are assumed to be bounded and positive definite as follows:
@@
We denote $d\mu=x_{d}^{\alpha}dxdt$, $d\mu=x_{d}^{\alpha}dx$, or $d\mu=x_{d}^{\alpha}dx_{d}$. It will always be clear from context which measure is used in a particular situation. We say that $u$ is a weak solution to (1.1)-(1.2) if
$\begin{gathered}\int_{\Omega_{T}}(-a_{0}u\cdot\varphi_{t}+(A^{ij}D_{j}u+x_{d}^{-1}B^{i}u)\cdot D_{i}\varphi+x_{d}^{-1}\hat{B}^{i}D_{i}u\cdot\varphi+x_{d}^{-2}Cu\cdot\varphi+\lambda C_{0}u\cdot\varphi)d\mu\\ =\int_{\Omega_{T}}(F^{i}\cdot D_{i}\varphi+x_{d}^{-1}F_{0}\cdot\varphi+\sqrt{\lambda}f\cdot\varphi)d\mu\end{gathered}$ (2.3)
holds for all $\varphi\in C_{0}^{\infty}((-\infty,T)\times\mathbb{R}^{d})^{n}$. For the integrals above to be well defined, we require
$x_{d}^{\alpha}(|u|+|Du|+x_{d}^{-1}|B||u|+x_{d}^{-1}|\hat{B}||Du|+x_{d}^{-2}|C||u|+|F|+x_{d}^{-1}|F_{0}|+|f|)\in L_{1,\text{loc}}((-\infty,T)\times\overline{\mathbb{R}^{d}_{+}}).$
We say that $u$ is a weak solution to (1.1) if (2.3) holds for all $\varphi\in C_{0}^{\infty}((-\infty,T)\times\mathbb{R}^{d}_{+})^{n}$.
For $p,q\in(1,\infty)$, $\infty\leq S<T\leq+\infty$, an open set $\mathcal{O}\subset\mathbb{R}^{d}_{+}$, and weights (that is, measurable and almost everywhere positive finite functions) $w_{0}(t)$ and $w_{1}(x)$, we define $L_{q,p}((S,T)\times\mathcal{O},wd\mu)$ to be the weighted mixed-norm Lebesgue space with the norm
$\|f\|_{L_{q,p}((S,T)\times\mathcal{O},wd\mu)}=\left(\int_{S}^{T}\left(\int_{\mathcal{O}}|f(t,x)|^{p}w_{1}(x)d\mu(x)\right)^{q/p}w_{0}(t)dt\right)^{1/q}.$
For $\mathcal{O}=\mathbb{R}^{d}_{+}$, we set $\mathcal{O}_{\text{test}}=\mathbb{R}^{d}$. For $\mathcal{O}=D_{r}^{+}(x_{0})$ with $x_{0}\in\overline{\mathbb{R}^{d}_{+}}$ and $x_{0d}\leq r$, we set $\mathcal{O}_{\text{test}}=B_{r}^{\prime}(x_{0}^{\prime})\times(-\infty,x_{0d}+r)$. For $\mathcal{O}=D_{r}(x_{0})$ with $x_{0}\in\overline{\mathbb{R}^{d}_{+}}$ and $x_{0d}>r$, we set $\mathcal{O}_{\text{test}}=D_{r}(x_{0})$. For such $\mathcal{O}$, we define
$\mathbb{H}^{-1}_{q,p}((S,T)\times\mathcal{O},wd\mu)=\big\{g\in\mathcal{D}^{\prime}((S,T)\times\mathcal{O}_{\text{test}}):\;\text{for some $F^{i},f\in L_{q,p}((S,T)\times\mathcal{O},wd\mu)$}$
$\text{and for all}\;\varphi\in C_{0}^{\infty}((S,T)\times\mathcal{O}_{\text{test}}),\;\langle g,\varphi\rangle=\int_{(S,T)\times\mathcal{O}}(F^{i}\cdot D_{i}\varphi+f\cdot\varphi)d\mu\big\},$ (2.4)
@@
$\mathcal{H}^{1}_{q,p}((S,T)\times\mathcal{O},wd\mu)=\{u:u,Du\in L_{q,p}((S,T)\times\mathcal{O},wd\mu),x_{d}^{\alpha}a_{0}(x_{d})u_{t}\in\mathbb{H}^{-1}_{q,p}((S,T)\times\mathcal{O},wd\mu)\},$
$\|u\|_{\mathcal{H}^{1}_{q,p}((S,T)\times\mathcal{O},wd\mu)}=\|Du\|_{L_{q,p}((S,T)\times\mathcal{O},wd\mu)}+\|u\|_{L_{q,p}((S,T)\times\mathcal{O},wd\mu)}+\|x_{d}^{\alpha}a_{0}(x_{d})u_{t}\|_{\mathbb{H}^{-1}_{q,p}((S,T)\times\mathcal{O},wd\mu)}.$
Here we understand $x_{d}^{\alpha}a_{0}(x_{d})u_{t}$ as an element of $\mathcal{D}^{\prime}((S,T)\times\mathcal{O}_{\text{test}})$ given by
@@
for $\varphi\in C_{0}^{\infty}((S,T)\times\mathcal{O}_{\text{test}})$. Note that the definitions of the spaces $\mathbb{H}^{-1}_{q,p}$ and $\mathcal{H}^{1}_{q,p}$ depend on the choice of the coefficient matrix $a_{0}(x_{d})$.
@@
For a set $\Omega$, a measure $\nu$ on $\Omega$ with $\nu(\Omega)<\infty$, and a function $f\in L_{1}(\Omega,\nu)$, we write
$\fint_{\Omega}fd\nu=\frac{1}{\nu(\Omega)}\int_{\Omega}fd\nu.$
@@
Similarly, a weight $w$ on $\mathbb{R}$ belongs to the class $A_{p}(\mathbb{R})$ if
$w]_{A_{p}(\mathbb{R})}=\sup_{r>0,t\in\mathbb{R}}\left(\fint_{t-r^{2}}^{t+r^{2}}w(s)ds\right)\times\left(\fint_{t-r^{2}}^{t+r^{2}}w^{-\frac{1}{p-1}}(s)ds\right)^{p-1}<\infty.$
@@
For compactness, from now on and throughout the paper the numbers $T,K$, and $K_{1}$ are considered given in the following ranges and used exclusively in the following contexts without further specifying.
  1. (1)
    $T\in(-\infty,\infty]$ is the terminal time.
  2. (2)
    All $A_{p},X_{p}$, and $Y_{p}$ constants are bounded by a constant $K\in(1,\infty)$. That is, whenever we write $w\in Z_{s}(\cdot)$ with $Z\in\{A,X,Y\}$ and $s\in(1,\infty)$, we assume $w]_{Z_{s}(\cdot)}\leq K$.
  3. (3)
    The constant $K_{1}\in[0,\infty)$ is involved in the growth bounds on the coefficients $B,\hat{B}$, and $C$.

### v2

@@ Abstract.
We prove the well-posedness and regularity of solutions in mixed-norm weighted Sobolev spaces for a class of second-order parabolic and elliptic systems in divergence form in the half-space $\mathbb{R}^{d}_{+}=\{x_{d}>0\}$ subject to the conormal boundary condition. Our work extends results previously available for scalar equations to the case of systems of equations. The leading coefficients are the product of $x_{d}^{\alpha}$ and bounded non-degenerate matrices, where $\alpha\in(-1,\infty)$. The leading coefficients are assumed to be merely measurable in the $x_{d}$ variable, and to have small mean oscillations in small cylinders with respect to the other variables. If the parameter $\alpha>0$, the lower-order coefficients are allowed to blow-up near the boundary. Our results readily generalize to infinite-dimensional equations in general real and complex Hilbert spaces.
@@ 1. Introduction
Here for $i,j=1,\dots,d$, the coefficients $A^{ij}$, $B^{i}$, $\hat{B}^{i}$, $C$, $C_{0}:\Omega_{T}\to\mathbb{R}^{n\times n}$ and $a_{0}=a_{0}(x_{d}):\mathbb{R}_{+}\to\mathbb{R}^{n\times n}$ are $n\times n$ matrix-valued measurable functions on $\Omega_{T}$, $F^{i},F_{0},f:\Omega_{T}\to\mathbb{R}^{n}$ are given (column-) vector-valued measurable functions on $\Omega_{T}$, and $u:\Omega_{T}\to\mathbb{R}^{n}$ is the unknown solution. That is, $u=[u^{1},\dots,u^{n}]^{T}\in\mathbb{R}^{n}$, and similarly for $F^{i},F_{0}$, and $f$; $A^{ij}=[A^{ij,k}_{l}]_{1\leq k,l\leq n}\in\mathbb{R}^{n\times n}$, and similarly for $a_{0},B^{i},\hat{B}^{i},C$, and $C_{0}$. We impose standard non-degeneracy, boundedness, and strong ellipticity conditions on $a_{0},C_{0}$, and $A^{ij}$; see (2.1) and (2.2) in Section 2. We naturally identify $n\times n$ matrices (e.g., the coefficients $a_{0},A^{ij},B^{i},\hat{B}^{i},C,C_{0}$) with linear operators $\mathbb{R}^{n}\to\mathbb{R}^{n}$. Here $Cu$ stands for the product of a matrix $C$ and a column-vector $u$, that is, $(Cu)^{k}=\sum_{l=1}^{n}C^{k}_{l}u^{l}$ for $1\leq k\leq n$, and similarly for the other terms in (1.1). Moreover, by $D_{i}=\frac{\partial}{\partial x_{i}}$ we denote the spatial weak partial derivatives, $\lambda\geq 0$ is a parameter, and we employ the summation convention over repeated upper and lower indices throughout the paper.
@@
We show that under some regularity assumptions on $A^{ij}$ and $C_{0}$, for any $p,q\in(1,\infty)$ and wide classes of weights $w=w(t,x)=w_{0}(t)w_{1}(x)$, holds the estimate
$\|Du\|_{L_{q,p}(\Omega_{T},wd\mu)}+\sqrt{\lambda}\|u\|_{L_{q,p}(\Omega_{T},wd\mu)}\leq N(\|F\|_{L_{q,p}(\Omega_{T},wd\mu)}+\|f\|_{L_{q,p}(\Omega_{T},wd\mu)}),$ (1.4)
where $N>0$ does not depend on $u,F$, $f$, and also the dimension $n$. Here $L_{q,p}(\Omega_{T},wd\mu)$ is a mixed-norm weighted Lebesgue space with the norm
$\|f\|_{L_{q,p}(\Omega_{T},wd\mu)}=\left(\int_{-\infty}^{T}\left(\int_{\mathbb{R}^{d}_{+}}|f(t,x)|^{p}w_{1}(x)x_{d}^{\alpha}dx\right)^{q/p}w_{0}(t)dt\right)^{1/q},$
$d\mu$ stands for $x_{d}^{\alpha}dxdt$, which is a natural measure for $L_{p}$ estimates for (1.1)-(1.2), and we assume that $\lambda>0$ is sufficiently large. In addition to estimate (1.4), we also obtain the corresponding existence and uniqueness of solutions. See Theorem 2.2 (i) and Section 2 for precise statements and definitions. In the case of scalar equations ($n=1$), the result of our Theorem 2.2 (i) was known from the recent work [18]. The arguments in [18] cannot be directly generalized to the case of systems due to the use of Moser’s iteration, a purely real scalar technique, in one of the key steps of the proof. Showing that estimate (1.4) and solvability of (1.1)-(1.2) still hold in the case of systems with $n\geq 2$ was one of our motivations to write this paper. Moreover, our results hold for solutions taking values in general real or complex Hilbert spaces $H$. That is, if $u,F^{i},F_{0},f:\Omega_{T}\to H$ are $H$-valued, and the coefficients $a_{0},A^{ij},B^{i},\hat{B}^{i},C,C_{0}:\Omega_{T}\to\mathcal{B}(H)$, where $\mathcal{B}(H)$ is the space of bounded linear operators on $H$. However, we impose more restricitive measurability conditions on the operator-valued coefficients if the space $H$ is non-separable, see Section 7 for details.
@@
In [17] and [18] the authors considered scalar ($n=1$) non-divergence form analogues of (1.1)-(1.2) in the ranges $\alpha\in(-1,1)$ and $\alpha\in(-1,\infty)$, respectively. There, an additional structural condition $A^{dj}/A^{dd}\equiv\text{const}$ for $1\leq j\leq d-1$ is imposed. The results in [17, 18] essentially use the scalar structure of the equations and are not readily generalized to systems of equations. Moreover, it remains an open problem whether classical $W^{2}_{p}$ estimates hold for uniformly elliptic and parabolic systems with partially VMO coefficients. We therefore do not consider systems in non-divergence form in this paper and leave this question for future investigation.
@@
Recent results on boundary Schauder-type estimates for parabolic equations with the conormal boundary condition can be found in [1, 2, 10], and for elliptic equations in [11, 49]. In these works, Schauder estimates for equations with singular-degenerate coefficients are used to derive boundary Harnack principles for equations with bounded uniformly elliptic coefficients. Parabolic and elliptic systems in divergence form with the zero Dirichlet boundary condition were studied in [15], whereas scalar equations in non-divergence form were considered in [19]. In the recent work [3], the authors studied the solvability of equations in divergence form with the inhomogeneous Dirichlet boundary data taken in spaces of distributions with negative order of differentiability. In these papers, the natural range of the parameter $\alpha$ for equations with the Dirichlet boundary condition is $\alpha\in(-\infty,1)$.
@@
Various phenomena in pure and applied mathematics are modeled by equations whose coefficients in front of $u_{t},D^{2}u,Du$, and $u$ may exhibit different rates of decay or blow-up. Examples of such anisotropy in coefficients include the Black-Scholes-Merton equation in mathematical finance, degenerate viscous Hamilton-Jacobi equations, and equations involving Baouendi-Grushin operators arising in sub-Riemannian geometry. For recent results on weighted $L_{p}$ estimates for such anisotropic equations we refer the reader to [24, 23, 21, 22, 46, 26].
@@
$L_{p}$ estimates for elliptic and parabolic equations with unbounded lower-order terms is an extensively studied topic, particularly in the classical case $\alpha=0$. Such results are often obtained under some integrability assumptions on the coefficients $B,\hat{B}$, and $C$; common conditions include $B,\hat{B}\in L_{d}$, $C\in L_{d/2}$, or their generalizations. Under these assumptions, $L_{p}$ estimates can be established for both Dirichlet and conormal boundary conditions. For detailed results in this direction, see [32, 42, 38] for elliptic equations and [37] for parabolic equations. As we mentioned earlier, $L_{p}$ estimates for equations with the zero Dirichlet boundary condition under assumptions of type (1.5) are also well studied in the literature; see [13, 34, 39].
Lastly, we note that the literature on infinite-dimensional parabolic equations in general Banach spaces, including Hilbert spaces, is highly active and extensive. We refer the reader to [27] for semilinear stochastic evolution equations in Hilbert spaces with applications to stochastic optimal control and Hamilton-Jacobi-Bellman equations, [45] for stochastic optimal control problems in Banach spaces; see also [50, 51] and references therein. For deterministic equations, numerous maximal $L^{p}-L^{q}$ regularity results are available for equations in the so called UMD Banach spaces, an example of which are Hilbert spaces. See [8] for general parabolic initial-boundary value problems, [43, 44] for generalizations to weighted spaces, and references therein. We note that parabolic equations in UMD spaces are usually studied via an approach quite different from the one in the present paper. There, parabolic equations are usually viewed as abstract evolution equations and studied via functional analytic tools such as the theory of sectorial operators and $H^{\infty}$-calculus.
@@
In our proofs, we do not rely on any specific properties of $\mathbb{R}^{n}$ other than its Hilbert space structure, nor do we use the base field being $\mathbb{R}$ in any essential way. Since our initial motivation was studying systems of equations, we chose to write the proofs in the language of systems. We briefly explain in Section 7 how our results generalize to equations in general real or complex Hilbert spaces. The Hilbert space structure, however, is essential for deriving the $L_{2}$ estimates in Section 3. It is of interest to understand whether our results can be generalized to equations in more general spaces, for example, UMD spaces.
The rest of the paper is organized as follows. In Section 2 we introduce the notation, function spaces, and state the main results, Theorems 2.2 and 2.3. In Section 3 we prove the existence and uniqueness of $L_{2}$ solutions. In Section 4 we consider systems having simple coefficients. In Section 5 we prove Part (i) of Theorem 2.2, where $B,\hat{B}$, and $C$ satisfy (1.3). Section 6 addresses systems with singular lower-order terms satisfying (1.5), completing the proofs of Theorems 2.2 and 2.3. Finally, in Section 7 we discuss the generalization of our results to Hilbert space-valued solutions.
@@ 2.1. Notation
We equip $\mathbb{R}^{n}$ with the standard dot product $a\cdot b:=\sum_{k=1}^{n}a^{k}b^{k}$.
@@
By $D_{i}u=\frac{\partial u}{\partial x_{i}}$, $i=1,\dots,d$, we denote the spatial weak partial derivatives of $u$, $Du=(D_{1}u,\dots,D_{d}u)$, and $D_{x^{\prime}}u=(D_{1}u,\dots,D_{d-1}u)$.
Throughout the paper, for $A=[A^{ij}]^{1\leq i,j\leq d}$, $A^{ij}\in\mathbb{R}^{n\times n}$, we denote
$|A|=\sup_{\xi,\eta\in(\mathbb{R}^{n})^{d},|\xi|=|\eta|=1}\left|\sum_{i,j=1}^{d}A^{ij}\xi_{j}\cdot\eta_{i}\right|,$
where for $\xi=(\xi_{1},\dots,\xi_{d})\in(\mathbb{R}^{n})^{d}$ we set $|\xi|^{2}=\sum\limits_{i=1}^{d}|\xi_{i}|^{2}$, and similarly for $\eta$. In particular, $|A^{ij}\xi_{j}\cdot\eta_{i}|\leq|A||\xi||\eta|$ for all $\xi,\eta\in(\mathbb{R}^{n})^{d}$. For a matrix $C\in\mathbb{R}^{n\times n}$, by $|C|$ we denote its operator norm
$|C|:=|C|_{\text{op}}=\sup_{u\in\mathbb{R}^{n},|u|=1}|Cu|.$
For $B=(B^{1},\dots,B^{d})$ and $\hat{B}=(\hat{B}^{1},\dots,\hat{B}^{d})$, where $B^{i},\hat{B}^{i}\in\mathbb{R}^{n\times n}$, we set
$|B|_{c}^{2}=\left|\sum_{i=1}^{d}(B^{i})^{T}B^{i}\right|_{\text{op}},\quad|\hat{B}|_{r}^{2}=\left|\sum_{i=1}^{d}\hat{B}^{i}(\hat{B}^{i})^{T}\right|_{\text{op}}.$
Here $|\cdot|_{c}$ stands for “column” operator norm and $|\cdot|_{r}$ stands for “row” operator norm. In particular, for all $u,\xi_{1},\dots,\xi_{d}\in\mathbb{R}^{n}$ it holds that
$|(B^{1}u,\dots,B^{d}u)|=\sqrt{\sum_{i=1}^{d}|B^{i}u|^{2}}\leq|B|_{c}|u|,\quad\left|\sum_{i=1}^{d}\hat{B}^{i}\xi_{i}\right|\leq|\hat{B}|_{r}|\xi|.$
[…]
@@
$\kappa\sum_{i=1}^{d}|\xi_{i}|^{2}\leq\sum_{i,j=1}^{d}A^{ij}\xi_{j}\cdot\xi_{i},\quad|A|\leq\kappa^{-1},$ (2.1)
on $\Omega_{T}$ for all $\xi_{i}\in\mathbb{R}^{n},i=1,\dots,d$. The coefficient matrices $a_{0}$ and $C_{0}$ are assumed to be bounded and positive definite as follows:
@@
Let $\mathcal{O}\subset\mathbb{R}^{d}_{+}$ denote an open set of the form $\mathcal{O}=\mathcal{O}^{\prime}\times(a,b)$, where $\mathcal{O}^{\prime}\subset\mathbb{R}^{d-1}$ is open and $0\leq a<b\leq\infty$. For such $\mathcal{O}$, we set $\mathcal{O}_{\text{test}}=\mathcal{O}^{\prime}\times(-\infty,b)$ if $a=0$, and $\mathcal{O}_{\text{test}}=\mathcal{O}$ if $a>0$. Also let $\infty\leq S<T\leq+\infty$.
We denote $d\mu=x_{d}^{\alpha}dxdt$, $d\mu=x_{d}^{\alpha}dx$, or $d\mu=x_{d}^{\alpha}dx_{d}$. It will always be clear from context which measure is used in a particular situation. We say that $u$ is a weak solution to equation (1.1) in $(S,T)\times\mathcal{O}$ if
$\begin{gathered}\int_{(S,T)\times\mathcal{O}}(-a_{0}u\cdot\varphi_{t}+(A^{ij}D_{j}u+x_{d}^{-1}B^{i}u)\cdot D_{i}\varphi+x_{d}^{-1}\hat{B}^{i}D_{i}u\cdot\varphi+x_{d}^{-2}Cu\cdot\varphi+\lambda C_{0}u\cdot\varphi)d\mu\\ =\int_{(S,T)\times\mathcal{O}}(F^{i}\cdot D_{i}\varphi+x_{d}^{-1}F_{0}\cdot\varphi+\sqrt{\lambda}f\cdot\varphi)d\mu\end{gathered}$ (2.3)
holds for all $\varphi\in C_{0}^{\infty}((S,T)\times\mathcal{O};\mathbb{R}^{n})$. If $a=0$, we say that $u$ is a weak solution to equation (1.1) in $(S,T)\times\mathcal{O}$ with the conormal boundary condition (1.2) satisfied on $(S,T)\times\mathcal{O}^{\prime}$ if (2.3) holds for all $\varphi\in C_{0}^{\infty}((S,T)\times\mathcal{O}_{\text{test}};\mathbb{R}^{n})$. For the integrals in (2.3) to be well defined, we will always have
$x_{d}^{\alpha}(|u|+|Du|+x_{d}^{-1}|B||u|+x_{d}^{-1}|\hat{B}||Du|+x_{d}^{-2}|C||u|+|F|+x_{d}^{-1}|F_{0}|+|f|)\in L_{1,\text{loc}}\left((S,T)\times\left(\mathcal{O}_{\text{test}}\cap\overline{\mathbb{R}^{d}_{+}}\right)\right).$
For $p,q\in(1,\infty)$, weights (that is, measurable and almost everywhere positive finite functions) $w_{0}(t)$ and $w_{1}(x)$, and a Banach space $X$, we define $L_{q,p}((S,T)\times\mathcal{O},wd\mu;X)$ to be the weighted mixed-norm Lebesgue space of measurable (strongly measurable in the sense of Bochner if $\operatorname{dim}X=\infty$) $X$-valued functions with the norm
$\|f\|_{L_{q,p}((S,T)\times\mathcal{O},wd\mu;X)}=\left(\int_{S}^{T}\left(\int_{\mathcal{O}}|f(t,x)|_{X}^{p}w_{1}(x)d\mu(x)\right)^{q/p}w_{0}(t)dt\right)^{1/q}.$
If $X=\mathbb{R}$ or $X=\mathbb{C}$, we write $L_{q,p}(\cdot)$ instead of $L_{q,p}(\cdot;X)$. When the space $X$ is understood from context, we will often write $f\in L_{q,p}(\cdot)$ meaning $|f|_{X}\in L_{q,p}(\cdot)$.
In general, if $(v_{k})$ is a finite collection of vectors $v_{k}\in\mathbb{R}^{n}$, we set $|(v_{k})|^{2}=\sum_{k}|v_{k}|^{2}$. For instance, $|Du|^{2}=\sum_{i=1}^{d}|D_{i}u|^{2}$, $|(F^{1},\dots,F^{d})|^{2}=\sum_{i=1}^{d}|F^{i}|^{2}$, etc.
We define
$\mathbb{H}^{-1}_{q,p}((S,T)\times\mathcal{O},wd\mu;\mathbb{R}^{n})=\big\{g\in\mathcal{D}^{\prime}((S,T)\times\mathcal{O}_{\text{test}};\mathbb{R}^{n}):\;\langle g,\varphi\rangle=\int_{(S,T)\times\mathcal{O}}(F^{i}\cdot D_{i}\varphi+f\cdot\varphi)d\mu$
$\text{for all}\;\varphi\in C_{0}^{\infty}((S,T)\times\mathcal{O}_{\text{test}};\mathbb{R}^{n})\;\text{for some $F^{1},\dots,F^{d},f\in L_{q,p}((S,T)\times\mathcal{O},wd\mu;\mathbb{R}^{n})$}\big\},$ (2.4)
@@
$\mathcal{H}^{1}_{q,p}((S,T)\times\mathcal{O},wd\mu;\mathbb{R}^{n})=\bigl\{u\in L_{q,p}((S,T)\times\mathcal{O},wd\mu;\mathbb{R}^{n}):$
$Du\in L_{q,p}((S,T)\times\mathcal{O},wd\mu),x_{d}^{\alpha}a_{0}(x_{d})u_{t}\in\mathbb{H}^{-1}_{q,p}((S,T)\times\mathcal{O},wd\mu;\mathbb{R}^{n})\bigr\},$
$\|u\|_{\mathcal{H}^{1}_{q,p}((S,T)\times\mathcal{O},wd\mu;\mathbb{R}^{n})}$
$=\|Du\|_{L_{q,p}((S,T)\times\mathcal{O},wd\mu)}+\|u\|_{L_{q,p}((S,T)\times\mathcal{O},wd\mu)}+\|x_{d}^{\alpha}a_{0}(x_{d})u_{t}\|_{\mathbb{H}^{-1}_{q,p}((S,T)\times\mathcal{O},wd\mu;\mathbb{R}^{n})}.$
Here we understand $x_{d}^{\alpha}a_{0}(x_{d})u_{t}$ as an element of $\mathcal{D}^{\prime}((S,T)\times\mathcal{O}_{\text{test}};\mathbb{R}^{n})$ given by
@@
for $\varphi\in C_{0}^{\infty}((S,T)\times\mathcal{O}_{\text{test}};\mathbb{R}^{n})$. Note that the definition of the space $\mathcal{H}^{1}_{q,p}$ depends on the choice of the coefficient matrix $a_{0}(x_{d})$.
@@
For a set $\Omega$, a measure $\nu$ on $\Omega$ with $0<\nu(\Omega)<\infty$, and a function $f\in L_{1}(\Omega,\nu)$, we write
$(f)_{\Omega,\nu}=\fint_{\Omega}fd\nu=\frac{1}{\nu(\Omega)}\int_{\Omega}fd\nu.$
@@
Similarly, a weight $w$ on $\mathbb{R}^{m}$ belongs to the class $A_{p}(\mathbb{R}^{m})$ if
$w]_{A_{p}(\mathbb{R}^{m})}=\sup_{r>0,y_{0}\in\mathbb{R}^{m}}\left(\fint_{B_{r}(y_{0})}w(y)dy\right)\times\left(\fint_{B_{r}(y_{0})}w^{-\frac{1}{p-1}}(y)dy\right)^{p-1}<\infty.$